Exercise 2.3 Sales personnel for X Company are required to submit weekly reports

listing customer contacts made during the week. A sample of 61 weekly contact

reports showed a mean of 22.4 customer contacts per week for the sales personnel.


1
Expert's answer
2021-06-25T07:12:35-0400

a)

The critical value for "\alpha=0.05" and "df=n-1=61-1=60" degrees of freedom is "t_c= 2.000298."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{ s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{ s}{\sqrt{n}})"

"=(22.4-2.000298\times\dfrac{ 5}{\sqrt{61}}, 22.4+2.000298\times\dfrac{ 5}{\sqrt{61}})"

"\approx(21.11944, 23.68056)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "21.11944<\mu<23.68056," which indicates that we are 95% confident that the true population mean "\mu"  is contained by the interval "(21.11944, 23.68056)."


b)

The critical value for "\alpha=0.10" and "df=n-1=61-1=60" degrees of freedom is "t_c= 1.670649."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{ s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{ s}{\sqrt{n}})"

"=(22.4-1.670649\times\dfrac{ 5}{\sqrt{61}}, 22.4+1.670649\times\dfrac{ 5}{\sqrt{61}})"

"\approx(21.33048, 23.46952)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "21.33048<\mu<23.46952," which indicates that we are 90% confident that the true population mean "\mu"  is contained by the interval "(21,33048, 23.46952)."


Level of significance is a statistical term for how willing you are to be wrong. With a 95 percent confidence interval, you have a 5 percent chance of being wrong. With a 90 percent confidence interval, you have a 10 percent chance of being wrong. A 99 percent confidence interval would be wider than a 95 percent confidence interval. A 90 percent confidence interval would be narrower.